### v1

@@ 1 Introduction
MR images acquired using different imaging sequences can provide complementary soft tissue contrasts, which are collectively used for clinical disease diagnosis. However, MRI is a slow imaging modality, resulting in a high sensitivity to subject motion. Long acquisition time is also undesirable, as it leads to lower patient throughput than other popular imaging modalities. Accelerated MRI acquisition and reconstruction are highly desirable and remain an active research topic [16, 6, 14, 10, 11, 12]. In the past years, deep learning methods [8] have shown great potential to enable rapid MRI. Deep learning approaches typically use task-specific deep networks to learn image features associated with incomplete k-space sampling and aim to remove image artifacts and noises caused by k-space undersampling during accelerated acquisition. While the conventional single-task models can work well when the training and testing data stem from the same data distribution (e.g., acquired using the same imaging sequence), they typically underperform when training and testing data differ substantially (e.g., acquired from different imaging sequences) [1, 15, 9]. Therefore, multi-task deep learning using methods to synergistically learn image features across multiple image datasets with different image contrasts is highly desirable for robust and high-efficient reconstruction.
@@
This paper proposes a meta-learning framework to learn image features from multiple MR image datasets. Meta-learning is a stacking ensemble learning [4] scheme that is considered a process of “learning-to-learn”, which can learn to improve the learning algorithm and parameter generalizability over multiple training episodes, thus enabling each task to learn better [7]. In this study, we developed a bi-level meta-learning reconstruction framework (i.e., including base-level and meta-level) to handle the learning of multiple image datasets. At the base-level, we introduce new deep networks (e.g., base-learners) by unrolling proximal gradient descent in both image and k-space domains to cross-learn the image and frequency domain features for single image contrast. In the meta-level, we introduce an optimization algorithm that can alternatively optimize the base-learners and one additional meta-learner for efficiently characterizing mutual correlation among multiple image datasets. Our bi-level meta-learning can simultaneously reconstruct highly-undersampled k-space data acquired using different imaging sequences with an optimal reconstruction for all image contrasts.
@@
1. A novel bi-level meta-learning framework is proposed to reconstruct highly-undersampled MRI datasets acquired using different imaging sequences.
@@
3. The proposed algorithm is validated for reconstructing a set of knee MRIs acquired at different imaging contrasts and planes.
@@ 2.1 Forward Learnable Optimization using Gradient Descent
We propose to develop an unrolled network to solve both the upper-level variational minimization problem and the lower-level constrained optimization problem in model (1). We introduce a proximal gradient descent [13] inspired algorithm 1, which can be parametrized as an iterative procedure with total $T$ iterations. The input $\mathbf{x}^{(0)}_{i,j}$ of the algorithm is obtained from an initial network that tries to pre-estimate the missing k-space data (similar to conventional parallel imaging methods such as GRAPPA [6]) therefore providing a good initial start for Algorithm 1. This helps reduce the number of iterations needed for convergence and decreases the overall computational complexity. The overall procedure consists of two for-loops: an outer and an inner loop, as follows:
@@ 3.2 Experimental Results
Fig. 2 compares the reconstructed images for Sag-T2 and Cor-T2 contrasts from different methods at AR=6. While ISTA-Net and pMRI-Net can largely remove the aliasing artifacts at this undersampling level, both methods yielded suboptimal reconstruction performance with noticeable image blurring in the reconstructed images. Our proposed STL can enable better reconstruction than ISTA-Net and pMRI-Net with improved image sharpness, due to its ability to cross-learn both image and k-space features, therefore better removing noises and artifacts and meanwhile preserving high-frequency image features. Our proposed MTML using meta-learning performs the best, outperforming all single-task learning methods. The pixel-wise error maps show that MTML has the smallest reconstruction errors compared to the fully-sampled reference, and the error is more homogeneous across the entire image domain for MTML. The zoom-in images further highlight the superb reconstruction performance of MTML where the tissue boundaries are better characterized at all tissue types (cartilage, meniscus, and muscle), and tissue texture, sharpness, and conspicuity are well-preserved. Similar comparisons of reconstructed images for the Sag-PD and Cor-PD sequences are shown in the Appendix, where MTML also consistently enables the best and most balanced performance on different types of contrast.

### v2

@@ 1 Introduction
MR images acquired using different imaging sequences can provide complementary soft tissue contrasts, which are collectively used for clinical disease diagnosis. However, MRI is a slow imaging modality, resulting in a high sensitivity to subject motion. Long acquisition time is also undesirable, as it leads to lower patient throughput than other popular imaging modalities. Accelerated MRI acquisition and reconstruction are highly desirable and remain an active research topic [25, 14, 23, 19, 20, 21, 9]. In the past years, deep learning methods [17] have shown great potential to enable rapid MRI. Deep learning approaches typically use task-specific deep networks to learn image features associated with incomplete k-space sampling and aim to remove image artifacts and noises caused by k-space undersampling during accelerated acquisition. While the conventional single-task models can work well when the training and testing data stem from the same data distribution (e.g., acquired using the same imaging sequence), they typically underperform when training and testing data differ substantially (e.g., acquired from different imaging sequences) [1, 24, 18]. Therefore, multi-task deep learning using methods to synergistically learn image features across multiple image datasets with different image contrasts is highly desirable for robust and high-efficient reconstruction.
MR images acquired using different imaging sequences can provide complementary soft tissue contrasts, which are collectively used for clinical disease diagnosis. However, MRI is a slow imaging modality due to its inherent sequential acquisition mode, resulting in a high sensitivity to subject motion. Long acquisition time is also undesirable, as it leads to lower patient throughput than other popular imaging modalities such as Computed Tomography (CT) and Ultrasound. Accelerated MRI acquisition and reconstruction are highly desirable and remain an active research topic in the MRI research community [25, 14, 23, 19, 20, 21, 5, 8]. In the past years, deep learning methods [17] have shown great potential to enable rapid MRI. Deep learning approaches typically use task-specific deep networks to learn image features associated with incomplete k-space sampling and aim to remove image artifacts and noises caused by k-space undersampling during accelerated acquisition. The trained models can then be applied to reconstruct newly undersampled data in the testing phase. While these single-task models can work well when the training and testing data stem from the same data distribution (e.g., acquired using the same imaging sequence), they typically underperform when training and testing data differ substantially (e.g., acquired from different imaging sequences) [1, 24, 18]. Therefore, multi-task deep learning using methods to synergistically learn image features across multiple image datasets with different image contrasts is highly desirable for robust and high-efficient reconstruction.
@@
This paper proposes a meta-learning framework to learn image features from multiple MR image datasets. Meta-learning is a stacking ensemble learning [11] scheme that is considered a process of “learning-to-learn”, which can learn to improve the learning algorithm and parameter generalizability over multiple training episodes, thus enabling each task to learn better [15]. In this study, we developed a bi-level meta-learning reconstruction framework (i.e., including base-level and meta-level) to handle the learning of multiple image datasets. At the base-level, we introduce new deep networks (e.g., base-learners) by unrolling proximal gradient descent in both image and k-space domains to cross-learn the image and frequency domain features for single image contrast. In the meta-level, we introduce an optimization algorithm that can alternatively optimize the base-learners and one additional meta-learner for efficiently characterizing mutual correlation among multiple image datasets. Our bi-level meta-learning can simultaneously reconstruct highly-undersampled k-space data acquired using different imaging sequences with an optimal reconstruction for all image contrasts. Meta
@@
1. A novel bi-level meta-learning framework is proposed to reconstruct highly-undersampled MRI datasets acquired using different imaging sequences. The trained meta-learning model can achieve improved reconstruction performance superior to standard deep learning methods, which can only perform well on single-task reconstruction.
@@
3. The proposed algorithm is validated for reconstructing a set of knee MRIs acquired at different imaging contrasts and planes. With the testing at various acceleration rates, the proposed meta-learning demonstrates superb reconstruction performance through faithfully removing image artifacts and noises, preserving truthful image textures and details, and maintaining high image sharpness and conspicuity.
@@ 2.1 Forward Learnable Optimization using Gradient Descent
We propose to develop an unrolled network to solve both the upper-level variational minimization problem and the lower-level constrained optimization problem in model (1). We introduce a proximal gradient descent [22] inspired algorithm 1, which can be parametrized as an iterative procedure with total $T$ iterations. Proximal gradient descent algorithm has been largely implemented in various applications for MRI reconstruction [3, 5, 2, 7, 6, 10, 28]. The input $\mathbf{x}^{(0)}_{i,j}$ of the algorithm is obtained from an initial network that tries to pre-estimate the missing k-space data (similar to conventional parallel imaging methods such as GRAPPA [14]) therefore providing a good initial start for Algorithm 1. This helps reduce the number of iterations needed for convergence and decreases the overall computational complexity. The overall procedure consists of two for-loops: an outer and an inner loop, as follows:
@@ 3.2 Experimental Results
Table 2: The hyper-parameter selection. (All the methods were implemented using PyTorch, and training and testing were conducted using two NVIDIA A100 GPUs.)
Initializer Xavier [13] Learning rate for $\Theta$ $\beta=1e-4$ $L$ 100 $K$ $10$
Optimizer ADAM [16] Learning rate for PD $\alpha_{i}=5e-4$ $T$ $5$ $r$ $5$
Training data 20 subjects Learning rate for T2 $\alpha_{i}=2e-4$ $\rho_{i}^{(0)}$ $0.5$ $\delta_{0}$ $0.5$
Testing data 5 subjects Kernel size $3\times 3\times 32$ $\lambda$ $41e-$ 4 $\mu$ $1$
Image size $300\times 300$ Slices 30 to 38 $m$ $4$ $c$ $18$
@@
Fig. 2 compares the reconstructed images for Sag-T2 and Cor-T2 contrasts from different methods at AR=6. While ISTA-Net and pMRI-Net can largely remove the aliasing artifacts at this undersampling level, both methods yielded suboptimal reconstruction performance with noticeable image blurring in the reconstructed images. Our proposed STL can enable better reconstruction than ISTA-Net and pMRI-Net with improved image sharpness, due to its ability to cross-learn both image and k-space features, therefore better removing noises and artifacts and meanwhile preserving high-frequency image features. Our proposed MTML using meta-learning performs the best, outperforming all single-task learning methods. The pixel-wise error maps show that MTML has the smallest reconstruction errors compared to the fully-sampled reference, and the error is more homogeneous across the entire image domain for MTML. The zoom-in images further highlight the superb reconstruction performance of MTML where the tissue boundaries are better characterized at all tissue types (cartilage, meniscus, and muscle), and tissue texture, sharpness, and conspicuity are well-preserved. Similar comparisons of reconstructed images for the Sag-PD and Cor-PD sequences are shown in the Table 2, where MTML also consistently enables the best and most balanced performance on different types of contrast.